OpenStudy (anonymous):

Help with 2 questions? Will return favor.

OpenStudy (anonymous):

.The ray bisects the angle DGH so DGI will equal IGH. x - 3 = 2x - 13 10 = x DGI = 10 - 3 = 7. IGH = 20 - 13 = 7. Hope I helped! (:

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Your welcome (:

OpenStudy (anonymous):

@Dobby1

OpenStudy (dobby1):

x should be -5

OpenStudy (anonymous):

Can you work it out?

OpenStudy (dobby1):

shoot i just messed up the question x = -10

OpenStudy (anonymous):

I gaved you the answer as I explained @supsavannah

OpenStudy (dobby1):

and you were right i keep messing up but i just fpoung that the anser to x is 10

OpenStudy (dobby1):

♥ωιnnιε.тнε.ρσσh♥ was right

OpenStudy (dobby1):

sorry i cant help with this with out graph paper and i dont have any

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

The midpoint of segment_UV is (5, –11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V. : Let the other end point be x2, y2 Find x2 \[\frac{ 3+x2 }{ 2}\] = 5 Multiply both sides by 2 3 + x2 = 10

OpenStudy (dobby1):

i'll try

OpenStudy (anonymous):

ok thanks.

OpenStudy (dobby1):

i belive but i do not know how to do the work i put it in a math site but i need to actually pay mony to see the steps

OpenStudy (dobby1):

7.3

OpenStudy (anonymous):

ok, thanks.

OpenStudy (dobby1):

thx

OpenStudy (anonymous):

Is this the correct work?